Find 3 consecutive integers in which the sum of 1st and 3rd is 84 more than one half of the 2nd

Answer by checkley79(3341)

YOU HAVE:
X the first integer
X+1 the second integer
x+2 the third integer
x+(x+2)=84+(x+1)/2
2x+2=84+(x+1)/2
2x-(x+1)/2=84-2 combine the 2 x terms.
(4x-x-1)/2=82
(3x-1)/2=82 cross multiply
3x-1=2*82
3x-1=164
3x=164+1
3x=165
x=165/3
x=55 the first integer.
55+1=56 the second integer.
55+2=57 the third integer.
Proof:
55+57=84+56/2
112=84+28
112=112